# FT2TF: First-Person Statement Text-To-Talking Face Generation

### v1

@@ 4.4.1 Quantitative Results
As shown in Table 1, FT2TF not only demonstrates the effectiveness compared to existing models but also achieves state-of-the-art performance across multiple evaluation metrics in talking face generation. This is evidenced by the quantitative results obtained on both LRS2 and LRS3 datasets.
Video Quality Enhancement. In terms of video quality evaluation metrics (PSNR, SSIM, and LPIPS), FT2TF reaches the state-of-the-art. Specifically, in comparison with other whole-face generation approaches on LRS2, it demonstrates significant improvements, with PSNR ($33.20$), SSIM ($0.9901$), and LPIPS ($0.0058$) showing an improvement of at least $92.5\%$, $78.0\%$, and $97.4\%$, respectively. FT2TF consistently outperforms lip-specific methods, even those using ground truth for the remaining parts of the face. Similar findings can be observed for the LRS3. These results show that FT2TF-generated videos exhibit the highest quality and are visually the closest to real-world ones.
Data Distribution Alignment. FID represents the alignment quality of data distribution, in which FT2TF achieves the optimal scores of $22.52/15.97$ on LRS2/3 datasets. The significant decrease of FID (e.g. at least $-70.6\%$ on LRS2 compared to the highest performing face-generation models) indicates the best-aligned data distribution between synthesized frames and real-world ones. This clearly demonstrates that our generated videos are the most statistically indistinguishable from the real ones.
Lip Synchronization and Identity Preservation. Although FT2TF is not exclusively designed for lip generation, our lip synchronization results (LipLMD) outperform all face-generation methods and remain competitive with the latest lip-specific approach (IP_LAP [46]). Since our model learns the entire face representation, it preserves facial expressions completely, capturing every detail. Consequently, we reach the best identity preservation as presented by the CSIM score.
Robustness Across Datasets. FT2TF shows robust performance across distinct and non-overlap datasets. Despite the variations between the two datasets outlined in Section 4.1, FT2TF consistently shows its stable and high-quality generation capabilities. This highlights the model’s adaptability and its ability to generalize across diverse video sources.
Efficiency Analysis. Figure 4 compares the total trainable parameters across various models, including both lip-specific and face-generation approaches. Specifically, when compared with the state-of-the-art methods that only learn lip-around features instead of the whole face (IP_LAP [46], Wav2Lip [28]), our model achieves a higher SSIM score while using a competitive number of trainable parameters. Moreover, our model outperforms other cutting-edge face-generation models significantly. Quantitatively, we achieve a $78.01\%$ improvement in SSIM with only $2.47\%$ more trainable parameters than MakeItTalk [49]. In comparison to other face-generation models, we present substantial advantages in terms of both trainable parameters and performance, with an average performance improvement of $122.88\%$ and an average parameter decrease of $68.94\%$.
@@
Figure 5: Qualitative comparison with state-of-the-art methods on LRS2 and LRS3. The three models on the left specialize in lip generation, whereas the others are designed to generate entire faces. Our model consistently generates the most detailed and accurate talking faces across diverse roles, genders, and ages under different lighting and head-poses conditions.
@@
4.4.2 Qualitative Results
We process all talking face frames using the same face detector [28], enabling a more detailed comparison of facial expressions. Figure 5 shows the comparison of generation quality between our model and relevant methods on LRS2 and LRS3. Our results stand out as the most similar to the ground truth across diverse conditions, including different character roles and identities, various lighting, and face angles. Benefiting from the Global Emotion Text Encoder, Linguistic Text Encoder, and Multi-scale Cross-Attention modules, our model accurately preserves facial expression details, such as teeth, face shape, head pose, skin color, and texture details. Additionally, we also demonstrate the results for multiple video frames, as shown in Figure 6. A temporal comparison with the state-of-the-art face generation method [49] demonstrates our model constructs the most accurate and consistent video frames across the time domain. Besides the visual demonstration, we also report quantitative results of temporal generation in Section 4.5.
4.5 User Study
We conduct a user study on videos synthesized by the state-of-the-art lip/face generation methods. 30 participants with diverse roles, genders, and ages are invited to watch 4 videos from each method and evaluate on a scale of 1 to 5 for four different aspects: temporal transition smoothness, lip synchronization accuracy, facial detail naturalness, and overall video quality. According to the Mean Opinion Scores (MOS) reported in Table 2, our model highly improves the generation quality compared to state-of-the-art lip/face generation methods from all evaluation domains.
Figure 7: Demonstration of text-driven talking face manipulation. Talking face frames of the left and right sides are synthesized in response to two different textual inputs utilizing the same visual source.
4.6 Text-Driven Face Manipulation Analysis
As one of the valuable real-world applications, we conduct experiments on talking face manipulation based on different text inputs. As shown in Figure 7, we observe two different outputs generated from distinct input texts correspondingly. The first input text, “IF YOU WERE WONDERFUL”, results in generating a detailed facial expression characterized by a smiling face, forehead wrinkles, enhanced cheekbones, and more open eyes. These details contribute to a brighter and more cheerful facial expressions that are consistent with the emotional characteristics conveyed by the input text, reflecting happiness and joy. Meanwhile, the second input text – “IF YOU WERE FEAR” generates faces with contrasting expressions, featuring a more serious and somber representation. Specifically, the forehead does not exhibit prominent wrinkles and the cheekbones appear less pronounced. Moreover, the mouth shape aligns with the intense emotion described in the text “FEAR”. The generated faces accurately reflect the emotion of fear, emphasizing the model’s capacity to adapt facial features in response to different emotional contexts. The reality of our model in achieving these remarkable results can be attributed to the integration of the Global Text Emotion Encoder and the Global Cross-Attention mechanism. The Global Text Emotion Encoder captures emotional semantics from the input text, which is then incorporated by the Global Cross-Attention module for the reflection on different regions of the generated faces. This experimental result showcases the model’s potential in real-world applications such as emotion-driven avatar generation, video dubbing, and virtual assistant interfaces.
Method PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ FIDnormal-$\downarrow$ CSIMnormal-$\uparrow$
@@ 4.6 Text-Driven Face Manipulation Analysis
4.7 Ablation Analysis
We conduct a comprehensive analysis of the ablation studies on the FT2TF to investigate the impact of key components, including the Cross-Attention mechanisms, loss terms, and different visual encoder backbones. The primary goal is to elucidate the model’s performance and the importance of these elements in generating high-quality face images from first-person text statements.
@@ 4.7 Ablation Analysis
Method PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ FIDnormal-$\downarrow$ CSIMnormal-$\uparrow$
@@
Role of Loss Terms. We delve into the role of specific loss terms, $L_{syn}$ and $L_{disc}$, as shown in Table 4. $L_{syn}$ focuses on improving the naturalness of lip movements and facial expression details, which leads to a significant gain in all the quantitative metrics. $L_{disc}$ guides the model to generate more realistic frames through binary predictions by the discriminator, contributing to accurate and expressive faces. It is also reflected obviously in the table.
Backbone PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ FIDnormal-$\downarrow$ CSIMnormal-$\uparrow$
@@
Encoder Backbones. We investigate the effectiveness of different backbone structures of the Visual Encoder, including ViT with different scales (ViT-B/L) and ResBlocks. As shown in Table 5, our model achieves stable and effective results with various encoder backbones, which proves the flexibility of our pipeline structure. Moreover, since ResBlocks retains information at different scales and inherently captures spatial hierarchies in the visual input, it results in a lead in PSNR, SSIM, and CSIM compared with ViT. Considering the parameter and data efficiency, we utilize ResBlocks as the Visual Encoder backbone.
@@ 5 Conclusion
This paper addresses the challenge of generating talking faces through a novel pipeline, FT2TF, which utilizes first-person text captions to replace traditional audio input. Our approach ensures the creation of faces with natural and detailed expressions. By incorporating two large language models, we extract emotional and linguistic features from the text. Moreover, the Multi-Scale Cross-Attention module effectively fuses visual and textual features, contributing to the synthesis of emotionally expressed talking faces. Extensive experiments on the LRS2 and LRS3 datasets underscore the effectiveness of FT2TF, demonstrating its capability to generate realistic, natural, and emotionally expressive talking face from first-person statement text.
[…]
Supplementary Material
@@ A Implementation Details
Video Preprocessing. The facial cropping method is from [28]. To ensure efficient training, we set the input frame size to 96×96.
Audio Preprocessing. Following the methodology presented in [28], we compute Mel-spectrograms from 16kHz audio data using a window size of 800 and a hop size of 200.
Visual Encoder. This module plays a pivotal role in capturing spatial information from the talking face video. It consists of 18 layers of 2D convolutional blocks, each accompanied by Batch Normalization and ReLU activation functions. Additionally, a subset of these modules incorporates residual blocks and skip connections, enhancing its ability to model complex visual features.
Global Emotion Text Encoder. The Global Emotion Text Encoder leverages a pre-trained Emoberta [19] model to encode the overarching emotional tones within the caption text. We import SentenceTransformer from the sentence_transformers library in Python and specifically use the tae898/emoberta-base model.
Linguistic Text Encoder. We adopt a pretrained GPT-Neo [10] model as our Linguistic Text Encoder. The encoder is set up using the transformers library in Python. We import AutoTokenizer and GPTNeoModel from transformers library. The tokenizer is initialized with AutoTokenizer.from_pretrained(”EleutherAI/gpt-neo-2.7B”), and the model is initialized with GPTNeoModel.from_pretrained(”EleutherAI/gpt-neo-2.7B”), respectively.
Visual Decoder. It is composed of six sets of 2D transpose convolution blocks, with each block consisting of a 2D transpose convolution layer and two layers of 2D residual convolution blocks. This architectural design enables the transformation of text descriptions into a coherent and expressive sequence of video frames.
Loss Weight Tuning. $\lambda_{1,2,3}$ are the weight of three losses, fine-tuned for optimal performance. During the initial 300 epochs, we assigned 0.7 to $\lambda_{1}$, 0.09 to $\lambda_{2}$, and 0.21 to $\lambda_{3}$. Subsequently, we adjusted these weights, setting $\lambda_{1}$ to 0.9, $\lambda_{2}$ to 0.03, and $\lambda_{3}$ to 0.07. Additionally, we explore the structural aspects of the Face Synthesizer and the Discriminator, highlighting their respective contributions and characteristics.
Figure S8: Questions in user study. We design 4 questions for each method mentioned in the user study, aiming to evaluate the models from multi-dimensions.
@@
Figure S9: Example screenshot of users’ responses summary of different methods. The results demonstrate that user evaluations are significantly more positive for FT2TF compared to other methods.
Figure S10: Generation samples using uncurated random samples on LRS2 [1] Test Set. Each pair of rows presents a single sample, with the top row of each pair showing the generated talking face frames and the bottom row showing the ground truth.
Figure S11: Generation samples using uncurated random samples on LRS3 [2] Test Set. Each pair of rows presents a single sample, with the top row of each pair showing the generated talking face frames and the bottom row showing the ground truth.
B Loss Structural Insights
The Discriminator is composed of 14 layers of fully convolutional modules, without the use of Batch Normalization layers or skip connections. It operates as a binary classifier, providing binary predictions ($O_{pred}$) based on whether the input is generated or Ground Truth. These predictions are guided by the binary labels ($Y_{disc}$). This binary classification loss, as detailed in $L_{disc}$, aids in ensuring the quality and authenticity of the synthesized frames.
In summary, FT2TF’s loss functions, coupled with the structural characteristics of the Face Synthesizer and the Discriminator, facilitate a comprehensive training process. These components work synergistically to improve the pixel-level fidelity, lip synchronization, and overall realism of the generated talking face videos, making FT2TF an effective solution for natural and expressive talking face synthesis.
C Description of Compared Methods
Below is a brief overview of the compared methods:
  • ATVGnet [5] generates the region around the lip (e.g. nose+lip) by constructing the face landmark from input audio signal and reference image.
  • Wav2Lip [28] constructs a lip masking and reconstruction strategy for lip generation, with the input as audio and video frames.
  • IP_LAP [46] is proposed recently and is regarded as the state-of-the-art method on audio-driven talking face generation. It leverages information from source audio, video, and landmarks through a two-stage framework.
  • EAMM [17] uses pose sequence, emotion source, and audio information to implement talking face synthesis and manipulation.
  • PC-AVS [47] focuses on pose-controllable talking head generation by devising a low-dimension pose code.
  • MakeItTalk [49] disentangles input audio into content and speaker domains, and predicts 3D landmarks to animate the facial expressions.
D Ablation Study on Text Encoders
Method PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ FIDnormal-$\downarrow$ CSIMnormal-$\uparrow$
swapping TE 24.24 0.9181 46.68 0.8655
FT2TF (Ours) 33.20 0.9901 22.52 0.9642
Table S6: Ablation study on Text Encoders. We conduct an ablation study on swapping two Text Encoders (TE) on LRS2.
In the ablation study outlined in Table S6, we investigate the effects of interchanging Text Encoders introduced in the Method Section. The Global Emotional Text Encoder, initially employing Emoberta [19], is replaced by a Linguistic Text Encoder (GPT-Neo [10]) to more thoroughly capture the global linguistic features in caption text. Conversely, the original Linguistic Text Encoder, utilizing GPT-Neo for linguistic encoding of text corresponding to generated talking face frames, is replaced with the Global Emotional Text Encoder (Emoberta) to solely extract emotional features in the remaining sentence, which corresponds to the talking face frames being generated.
This Text Encoder swapping results in a notable decrease in performance metrics. The GPT-Neo exhibits limited capabilities in conveying emotional expressions or depicting facial details. In contrast, the Emoberta is confined to processing only parts of the sentence corresponding to the generated talking face frames, thus restricting its effectiveness in accurately capturing correct sentence-level emotional nuances. This shortcoming can lead to the extraction of inappropriate emotional cues. As a result, the model faces challenges in synchronizing facial details with the emotional aspects embedded in the ground truth during training, causing sub-optimal performance.
E User Study
@@
The user study questionnaire is shown in Figure S8. We design 4 questions for each method to let the participant answer. The questions include 4 different aspects: temporal transition smoothness, lip synchronization accuracy, facial detail naturalness, and overall video quality.
Response From Users.
Part of the responses from users are shown in Figure S9. From the results, we can observe that the users respond with much more positive results of FT2TF compared with other methods.
F Datasets
  • LRS2 [1]: The dataset consists of 48,164 video clips collected from BBC outdoor broadcasts, and is split into the train, validation, and test sets containing 45,839, 1,082, and 1,243 video clips, respectively. Each clip is paired with a text caption containing up to 100 characters. To ensure uniform sentence lengths during model training, we selectively retained videos from the training and test sets that had a total frame count falling within the range of 30 to 35 frames. Consequently, our training set comprises 6,406 videos, while the test set consists of 255 videos.
  • LRS3 [2]: The LRS3 dataset is constructed from over 400 hours of 5,594 indoor TED/TEDx shows with 151,819 videos, along with the corresponding text captions. Specifically, it consists of 32,000 train/val utterances and 1,452 test ones. Similarly, in line with the LRS2 dataset, we selected videos from the training and test sets within the 30 to 35 frame range. This resulted in a training set comprising 2,262 videos and a test set of 172 videos.
G Side-by-Side Comparison
We present more results of generated talking face video frames using uncurated random selections on the test sets of the LRS2 [1] and LRS3 [2] datasets. These results are displayed alongside the original video sequences for comparison, as shown in Figure S10 and Figure S11. The results showcase that our model can consistently generate detailed and accurate talking faces across diverse roles, genders, and ages under different lighting and head-poses conditions.

### v2

@@ 4.4.1 Quantitative Results
As shown in Table 1, FT2TF achieves state-of-the-art performance across multiple evaluation metrics for talking face generation, as evidenced by the quantitative results on both LRS2 and LRS3 datasets.
Video Quality Enhancement. FT2TF achieves state-of-the-art results across video quality evaluation metrics (PSNR, SSIM, and LPIPS). For example, on the LRS2 dataset, FT2TF outperforms TTFS [16], the current state-of-the-art T+V method, with a $9.6\%$ improvement in PSNR, a $4.9\%$ improvement in SSIM, and a significantly lower LPIPS value, reduced by $66.5\%$. Moreover, FT2TF consistently outperforms lip-specific methods, even those using ground truth for the remaining parts of the face, further validating its superior performance. Similar results are observed on the LRS3 dataset, highlighting FT2TF’s ability to generate talking faces with high-quality and visual realism, closely resembling real-world data.
Data Distribution Alignment. In terms of FID, which evaluates the alignment of data distributions, FT2TF achieves scores of $22.52$ on LRS2 and $15.97$ on LRS3, demonstrating competitive performance compared to state-of-the-art methods [16]. The low FID values indicate that FT2TF generates frames with distributions statistically close to real ones, underscoring the model’s capacity to produce natural-looking talking faces.
@@
Lip Synchronization and Identity Preservation. Although FT2TF is not specifically designed for lip generation, it achieves comparable lip synchronization performance (LipLMD) to the state-of-the-art lip-specific method, IP_LAP [49], and surpasses all whole-face generation methods. Furthermore, FT2TF excels in identity preservation, as evidenced by the highest CSIM scores. This underscores its capability to maintain accurate facial features and expressions across frames, ensuring the realistic and coherent generation of talking faces.
Robustness Across Datasets. FT2TF demonstrates consistent and robust performance across distinct and non-overlapping datasets. Despite the variations between the two datasets described in Section 4.1, FT2TF maintains stable, high-quality generation capabilities. This underscores the model’s adaptability and its strong generalization ability across diverse video sources.
Figure 5: Qualitative comparison with state-of-the-art methods on LRS2 and LRS3. The three models on the left specialize in lip generation, whereas the others are designed to generate entire faces. TTFS [16] and our model are text-driven, whereas the remaining methods are audio-driven. Our model consistently generates the most detailed and accurate talking faces across diverse roles, genders, and ages under different lighting and head-poses conditions.
Efficiency Analysis. Figure 4 compares the total trainable parameters across various models, including both lip-specific and face-generation approaches. Specifically, compared to the state-of-the-art methods that only learn lip-around features instead of the whole face (IP_LAP [49], Wav2Lip [31]), our model achieves a higher SSIM score while using a competitive number of trainable parameters. Moreover, our model outperforms other cutting-edge face generation models significantly. Quantitatively, we achieve a $78.01\%$ improvement in SSIM with only $2.47\%$ more trainable parameters than MakeItTalk [52]. In comparison to other face-generation models listed in the figure, we present substantial advantages in terms of both trainable parameters and performance, with an average performance improvement of $122.88\%$ and an average parameter decrease of $68.94\%$.
4.4.2 Qualitative Results
We process all talking face frames using the same face detector [31], enabling a more detailed comparison of facial expressions. Figure 5 shows the comparison of generation quality between our model and the relevant methods on LRS2 and LRS3. Our results stand out as the most similar to ground truth in diverse conditions, including different character roles and identities, various lighting, and face angles. Benefiting from the Global Emotion Text Encoder, Linguistic Text Encoder, and Multi-Scale Cross-Attention modules, our model accurately preserves facial expression details, such as teeth, face shape, head pose, skin color, and texture details. Furthermore, we evaluate multiple consecutive video frames, as shown in Figure 6. Specifically, a temporal comparison with the state-of-the-art face generation methods shows that our model constructs the most accurate and consistent video frames in the time domain. In addition to the visual demonstration, we also report quantitative results of temporal generation in Section 4.5.
4.5 User Study
We conduct a user study to evaluate videos synthesized by state-of-the-art lip and face generation methods. 30 participants with diverse roles, genders, and ages are invited to watch 4 videos from each method and evaluate on a scale of 1 to 5 for four different aspects: temporal transition smoothness, lip synchronization accuracy, facial detail naturalness, and overall video quality. According to the Mean Opinion Scores (MOS) reported in Table 2, our model outperforms the state-of-the-art lip/face generation methods from all evaluation domains.
@@ 4.5 User Study
4.6 Ablation Analysis
We conduct a comprehensive analysis of the ablation studies on the FT2TF to investigate the impact of key components, including the Cross-Attention mechanisms, loss terms, and different visual encoder backbones. The primary goal is to analyze the model’s performance and demonstrate the contribution of each component in generating high-quality face frames from first-person text statements.
[…]
@@ 4.6 Ablation Analysis
[…]
@@
Role of Loss Terms. We delve into the role of specific loss terms, $L_{syn}$ and $L_{disc}$, as shown in Table 4. $L_{syn}$ focuses on improving the naturalness of lip movements and facial expression details, which leads to a significant gain in all quantitative metrics. $L_{disc}$ guides the model to generate more realistic frames through binary predictions by the discriminator, contributing to accurate and expressive faces.
[…]
@@
Encoder Backbones. We investigate the effectiveness of different backbone structures of the Visual Encoder, including ViT with different scales (ViT-B/L) and ResBlocks. As shown in Table 5, our model achieves stable and effective results with various encoder backbones, demonstrating the flexibility of our pipeline structure. Moreover, since ResBlocks retains information at different scales and inherently captures spatial hierarchies in the visual input, it results in a lead in PSNR, SSIM, and CSIM compared to ViT. Considering the parameter and data efficiency, we utilize ResBlocks as the Visual Encoder backbone.
@@ 5 Conclusion
This paper addresses the challenge of generating talking faces through the FT2TF pipeline, which uses first-person text captions to replace traditional audio input. Our approach ensures the generation of faces with natural and detailed expressions. By incorporating two large language models, we extract emotional and linguistic features from the text. Moreover, the Multi-Scale Cross-Attention module effectively fuses visual and textual features, contributing to the synthesis of emotionally expressed talking faces. Extensive experiments on the LRS2 and LRS3 datasets highlight FT2TF’s effectiveness in generating realistic, natural, and emotionally expressive talking faces from first-person text.
@@ A Implementation Details
A.1 Video Preprocessing
For video data preprocessing, we employ the facial cropping method introduced in [31]. To further enhance computational efficiency, we standardize the input frame resolution to 96×96 pixels, balancing detail retention with processing speed.
A.2 Audio Preprocessing
In alignment with [31], we generate Mel-spectrograms from audio samples recorded at 16kHz. This approach involves setting a window size of 800 samples and a hop size of 200, ensuring temporal resolution appropriate for capturing the nuances of lip movements.
A.3 Visual Encoder
The Visual Encoder plays a critical role in modeling spatial features from the talking face video. It consists of a deep convolutional architecture with 18 2D convolutional layers, each accompanied by Batch Normalization and ReLU activation functions. In addition, a subset of these modules incorporates residual blocks and skip connections, enhancing its ability to model complex visual features.
A.4 Global Emotion Text Encoder
The Global Emotion Text Encoder leverages a pretrained Emoberta [20] model to encode the overarching emotional tones within the caption text. We import SentenceTransformer from the sentence_transformers library in Python and specifically use the tae898/emoberta-base model.
A.5 Linguistic Text Encoder
We adopt a pretrained GPT-Neo [10] model as our Linguistic Text Encoder. The encoder is set up using the transformers library in Python. We import AutoTokenizer and GPTNeoModel from transformers library. The tokenizer is initialized with AutoTokenizer.from_pretrained("EleutherAI/gpt-neo-2.7B"), and the model is initialized with GPTNeoModel.from_pretrained("EleutherAI/gpt-neo-2.7B"), respectively.
A.6 Visual Decoder
The Visual Decoder is composed of six sets of 2D transpose convolution blocks, with each block consisting of a 2D transpose convolution layer and two layers of 2D residual convolution blocks. This architectural design enables the transformation of text descriptions into a coherent and expressive sequence of video frames.
A.7 Loss Weight Tuning
$\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are the weight of three losses, fine-tuned for optimal performance. During the initial 300 epochs, we assign 0.7 to $\lambda_{1}$, 0.09 to $\lambda_{2}$, and 0.21 to $\lambda_{3}$. Afterward, we adjust these weights, setting $\lambda_{1}$ to 0.9, $\lambda_{2}$ to 0.03, and $\lambda_{3}$ to 0.07.
B Loss Structural Insights
We explore the structural aspects of the Face Synthesizer and the Discriminator, highlighting their respective contributions and characteristics.
The Discriminator is composed of 14 layers of fully convolutional modules, without the use of Batch Normalization layers or skip connections. It operates as a binary classifier, providing binary predictions ($O_{pred}$) based on whether the input is generated or Ground Truth. These predictions are guided by the binary labels ($Y_{disc}$). This binary classification loss, as detailed in $L_{disc}$, helps to ensure the quality and authenticity of the synthesized frames.
In summary, FT2TF’s loss functions, coupled with the structural characteristics of the Face Synthesizer and the Discriminator, facilitate a comprehensive training process. These components work in harmony to improve the pixel-level fidelity, lip synchronization, and overall realism of the generated talking face videos, making FT2TF an effective solution for natural and expressive talking face synthesis.
Figure S7: User study questions. We design four questions for each method in the user study to evaluate the quality of generated talking faces across multiple dimensions.
@@ B Loss Structural Insights
Figure S8: Summary of user responses across methods. The results indicate that user evaluations are notably more favorable for FT2TF compared to other methods.
C User Study
To evaluate the quality of FT2TF’s generated talking faces in comparison to existing methods, we conduct a comprehensive user study assessing participant impressions across multiple dimensions.
@@ C User Study
The questionnaire used in the user study is shown in Figure S7. Four questions are designed for each method, covering key aspects: temporal transition smoothness, lip synchronization accuracy, facial detail naturalness, and overall video quality.
User Responses.
Figure S8 provides a summary of user responses. The results indicate that participants responded significantly more positively to FT2TF compared to other methods.